OpenStudy (anonymous):

Find a relationship between x and y such that (x, y) is equidistant (the same distance) from the two points. (6, 3), (5, −9)

OpenStudy (anonymous):

The formula should be: \[( \frac{(x1 + x2)}{2} , \frac{(y1 + y2)}{2} ) \] Let (x1, y1) = (6, 3), and (x2, y2) = (5, −9) Now substitute this in above formula to find the equidistant point.